# Feature Identification via the Empirical NTK

### v2

@@ 4 Results for modular arithmetic (grokking)
Next, we study the alignment of top eNTK eigenvectors with the ground-truth Fourier features in a MLP trained on modular arithmetic (Gromov (2023)).
@@ 4.1 Review of the setup
In the experiments below, we report results for $p=29$, $n=512$ and $\alpha=0.7$. However, we find qualitatively similar results for different values of the hyperparameters (including e.g. the narrower model with $n=64$, 3 other values of $p$, and 5 different seeds).
With these hyperparameters, the model exhibits grokking (Power et al. (2022), Nanda et al. (2023)). Namely, some time after reaching 100% accuracy on the training set, it suddenly goes from 0% to 100% accuracy on the test set (see the left panel of Fig. 3), suggesting that it switched to using a more generalizable representation of the underlying modular arithmetic algorithm.
Moreover, in this case we know the exact solution that the model learns after grokking (Gromov (2023)). At the end of training, the model solves modular addition by learning the weights
@@ 4.2 Results
Figure 2: Full and layerwise eNTK spectrum of the modular arithmetic model after training to convergence (500 epochs). The left column displays the spectrum of the full eNTK, the middle column the spectrum of the layerwise eNTK wrt the parameters (9) in layer 1, and the right column the spectrum of the layerwise eNTK wrt the parameters (10) in layer 2.
@@
Time evolution of the spectrum. The second cliff is absent before grokking and appears at the grokking phase transition (see the right panel of Fig. 3). This suggests that feature-finding aside, perhaps the eNTK can be used to find phase transitions. The eigenvectors evolve smoothly on either side of the phase transition even as the eigenvalues jump. We comment on the evolution of the eigenvectors in Appendix D.
On the other hand, the first cliff is present already at initialization (see the r.h.s. of Fig. 3). This follows from the structure of the Layer 1 kernel, as we explain briefly in Appendix B, and suggests (in combination with the subsection below) that this model's use of Fourier features could perhaps have been predicted at initialization from the eNTK.
Figure 3: Spectral change at the grokking phase transition. Left: Train and test accuracy shows evidence for a grokking phase transition (sudden onset of generalization) around epoch 90. Right: at the same training time, the eNTK spectrum develops a kink at the base of the second cliff.
@@
We would like to understand if a rotation inside the first cliff eigenspace could surface the Fourier modes. Since the space of possible inputs forms a $p\times p$ toroidal lattice, we can test for this by ordering directions in the first cliff eigenspace by the magnitude of sum-of-squared differences between the images of adjacent lattice points. If the Layer 1 cliff was in fact a rotation of the Fourier modes, this way of picking a basis should separate low-frequency Fourier modes from high-frequency ones.
In practice, we apply the two-stage ``per-axis graph smoothness" algorithm described in Appendix C. Upon doing so, we find cleanly that the Layer 1 cliff does in fact contain directions aligned with the $a$ and $b$ Fourier families. This is shown on the right-hand side of Fig. 4.
Figure 4: Results from applying the two-stage per-axis graph smoothness algorithm described in Appendix C to rotate the eNTK eigenvector basis in the Layer 1 cliff. Left: heatmap showing the squared norm of the inner product of eNTK eigenvectors in the cliff with the $(\cos a,\sin a)$ or $(\cos b,\sin b)$ feature families. Right: heatmap showing the squared norm of the inner product of basis vectors in the Layer 1 cliff found from applying the two-stage graph smoothness algorithm described in Appendix C with the $(\cos a,\sin a)$ or $(\cos b,\sin b)$ feature families.
A similar procedure shows that the second cliff appearing in the eNTK spectrum after grokking contains the $a\pm b$ Fourier families. We discuss this in Appendix D.
@@ 5 Discussion
To summarize, we have provided evidence that eigenanalysis of the eNTK can surface learned features across toy benchmarks commonly used by the mechanistic interpretability community.
There are several potentially interesting directions for follow-up work.
At the most basic level, it could be interesting to apply the pipeline presented here to other MLPs that model different aspects of realistic data. For example, we could apply our method to extensions of TMS that model composed or hierarchical features, or to toy models of feature splitting, addressing a known weakness of SAEs.
Ultimately, we would like to make contact with realistic datasets/models. For this, it will be important to extend our pipeline to Transformers. As a natural first exercise we plan to apply the pipeline to a Transformer trained on modular arithmetic, but we would also like to study how our method interacts with Transformers trained on algorithms that require nontrivial attention (e.g. induction heads (Olsson et al. (2022))), and on larger models (e.g. Bricken et al. (2023), OthelloGPT (Li et al. (2023)) and others). The latter may require advances in the underlying implementation to let us approximate the eNTK at scale.
For the 1L models studied in this paper, it's easy to write down closed-form expressions for the kernels in terms of the weights and data vectors. (See Appendix B.) It could be interesting to understand analytically what properties of said weights and data were needed to get the clean mechanistic signals presented here.
Finally, our route to recovering the Fourier features in Section 4 required imposing a geometric prior of smoothness on the input space by hand.  [^6] Though not full a priori knowledge of the Fourier features. It would be nice to be able to identify features without using any structural prior beyond the raw data. To this end, perhaps we could learn such a prior from the kernel itself in a way that could generalize to other situations.  [^7] For example, as a naive first idea, perhaps we could replace the algorithm in Appendix C with one that constructs a graph designating inputs to the model as “neighbors” iff their eNTK rows look similar, then runs a Laplacian rotation on the learned graph to sort the eNTK eigenvectors. However, it’s not clear whether and when this idea would generalize in an interesting way to other situations.
We hope to say more about these directions in the future.
Acknowledgements
I am grateful to Ari Brill, Lauren Greenspan, Andrew Mack, Logan Smith, Lucas Teixeira and Dmitry Vaintrob for discussions and comments on the manuscript. This work was supported by PIBBSS and the Long-Term Future Fund.
@@ Appendix A Review of NTK theory
In this section, we review the derivation of Eq. (1) since it motivates this entire line of study. As mentioned in the introduction, a main claim of the NTK theory is that a neural network trained under gradient flow with MSE loss learns the function (1), (2) at the end of training provided that its trajectory stays in a ``lazy limit" where the drift in parameters throughout training is parametrically small.
The idea behind the derivation is that if we take a limit where the changes in a model's weights $W_{\mu}$ over training are small enough that we can truncate a Taylor expansion of the model in its weights to first order while maintaining a good approximation, then perhaps we can resum the linearized approximation to how the model changes at each step of gradient descent to get a closed-form formula for the function learned by the model.
@@
[…]
$f_{i}(x,W_{\mu}(t+1))-f_{i}(x,W_{\mu}(t))=\left.-\eta\sum_{j}\sum_{\alpha\in\mathcal{D}}\frac{d\mathcal{L}}{df_{j}(x_{\alpha})}\left[\sum_{\mu}\frac{df_{i}(x)}{dW_{\mu}}\frac{df_{j}(x_{\alpha})}{dW_{\mu}}\right]\right|_{W_{\mu}=W_{\mu}(t)}+\dots\,$, (13)
where the term in brackets on the RHS is the NTK, (2).  [^8] Throughout this section we also use that the NTK itself is time independent in the lazy limit.
Now suppose we're allowed to truncate $\dots$ while maintaining a controlled approximation to the function learned by the neural network. Moreover, let's specialize to MSE loss. This yields
@@
Appendix B Kernels used in this paper
@@
For the modular arithmetic experiment, $f(x)=W^{(2)}(W^{(1)}x)^{2}$ (6) implies the layer-1 Jacobian
@@
It could be interesting, and probably is tractable, to understand which properties of the weights and data lead to the mechanistic structure found in the main text. In general, I will leave this for future work.
As an appetizer however, note that an especially striking result in the main text - the fact that the Layer 1 eigenvalue cliff in the modular arithmetic problem was present already at initialization, and hence that the use of Fourier features for modular arithmetic could perhaps have been predicted at initialization with the eNTK - is easily explained by the structure of Eq. (18). Eq. (18) contains a parameter-agnostic factor of $x_{1}\cdot x_{2}$ which is zero for one-hot-encoded data points $(a,b)$ and $(a^{\prime},b^{\prime})$ unless either $a=a^{\prime}$ or $b=b^{\prime}$. This makes the Layer 1 kernel have nonzero values along each fixed-$a$ and fixed-$b$ line, and zeroes elsewhere. Moreover, the nonzero kernel entries at initialization are approximately identical for each choice of $a$ or $b$ by symmetry. Hence, the kernel at initialization has an approximate shift symmetry along the $a$ and $b$ axes, so is approximately diagonalized by Fourier waves on each axis, with (nonzero) eigenvalues related to the weight initializations.
In this example, the $x_{1}\cdot x_{2}$ factor in (18) relied on the model architecture (specifically, the fact that each one-hot-encoded data point only touches two columns of the Layer 1 weight matrix), and its conversion to a shift-symmetric kernel relied on the shift symmetry of the dataset, illustrating that the NTK (and its eigenstructure) inherits information about both already at initialization.
Appendix C Graph smoothness algorithm
To rotate the eigenvectors in the Layer 1 cliff and obtain the plot on the RHS of Fig. 4 in section 4, we apply the following two-stage ``per-axis graph smoothness" algorithm:
@@ Appendix C Graph smoothness algorithm
    penalizing cases where entries in the (length $p^{2}$) vector $v$ corresponding to lattice points that differ by 1 in the $a,b$ directions respectively are not the same. In practice, the matrices $L_{a}$, $L_{b}$ that realize (21), (22) are $L_{1}\otimes I_{p}$ and $I_{p}\otimes L_{1}$ respectively, for $L_{1}$ the dimension $(p,p)$ matrix with 2 on the diagonal and -1 along diagonals offset by 1 and at the wrap-around entries $(1,p)$ and $(p,1)$.
@@
    We diagonalize one of the two Laplacians, WLOG $L_{a}$, inside the span of the Layer 1 cliff. Concretely, we first compress $L_{a}$ by computing $A=c^{T}L_{a}c$ for $c\in\mathbb{R}^{N,k}$ the column-orthonormalized collection of eNTK eigenvectors in the Layer 1 cliff, then eigendecompose the compressed Laplacian finding $A=U\Sigma U^{T}$ with $U\in\mathbb{R}^{k,k}$, and finally rotate the eNTK cliff by the eigenvector matrix of the compressed Laplacian, $c^{\prime}=cU$. This sorts the directions in the cliff by $a$-smoothness.  [^9] In a sentence, this is because directions where we minimize $v^{T}L_{a}v$ in a given vector space are the smoothest directions in the vector space by the operational definition (21) of the Laplacian, and we can minimize this from an eigendecomposition by the Rayleigh-Ritz theorem.
@@
    We then take the directions in the cliff corresponding to the bottom half of the $L_{a}$ eigenvalues  [^10] I.e. the bottom $2\lfloor\frac{p}{2}\rfloor$ directions in practice. and diagonalize $L_{b}$ inside that span using the same procedure as step 2, sorting those directions by $b$-smoothness.
The logic is that if the Layer 1 eNTK cliff was spanned exactly by $(\cos a,\sin a),(\cos b,\sin b)$ Fourier families, then we would expect the spectrum wrt $L_{a}$ to contain degenerate near-zero values corresponding to the `$b$' Fourier modes that are constant in the `$a$' direction, followed by a tower of increasing pairs of nonzero eigenvalues corresponding to the sorted `$a$' Fourier modes. So the two-stage process would sequentially disentangle them, and indeed this is what we find.
Appendix D Evolution of the sum/difference Fourier modes throughout training in the modular arithmetic experiment
In this section, we show that in the modular arithmetic experiment,
  1. 1.
    The second eigenvalue cliff in the full eNTK spectrum (left-hand column of Fig. 2) approximately contains the `sum' and `difference' Fourier feature families $(\cos a\pm b,\sin a\pm b)$.
  2. 2.
    The features themselves form continuously to either side of the grokking phase transition.
To see this, we apply the graph smoothness algorithm in Appendix C to the sum and difference families, first constructing Laplacians
@@
that penalize variation along the diagonals, then diagonalizing first the sum then the difference Laplacian within the span of the second eigenvalue cliff. This situation is equivalent to two-stage diagonalization in the $a$ and $b$ directions up to a 45-degree rotation, so the logic of Appendix C goes through. In particular, $(a,b)$ and $(a+1,b+1)$ have the same aba-b, so the difference family is flat wrt $L_{\rm sum}$, and vice versa.
@@
Figure 5: Results from applying the two-stage graph smoothness algorithm described in Appendix C wrt the sum and difference Laplacians (23), (24) to rotate the eNTK eigenvectors in the second eigenvalue cliff of the modular arithmetic experiment at different epochs. Each plot is a heatmap showing the squared norm of the inner product of rotated basis vectors in the second eigenvalue cliff with the $(\cos a+b,\sin a+b)$ or $(\cos a-b,\sin a-b)$ feature families.
@@
The results are shown in Fig. 5. From the bottom right plot showing the results of disentanglement at the end of training, we see that the second eigenvalue cliff approximately contains the sum and difference families. (The heatmaps for each family form a ``>>" shape because stepping across the diagonal advances the argument of each Fourier mode by $2\cdot 2\pi k/p$, so the Laplacian energy is minimal near $k\approx 0$ and $k\approx p/2$, and maximal near $k\approx p/4$.) The disentanglement is imperfect because the low-energy k=1``k=1" and k=14``k=14" sum modes mix with the $L_{\rm diff}$ modes  [^11] Indeed, we can make the heatmaps in Fig 5 appear closer to 1:1 by diagonalizing $L_{\rm diff}$ in Step 3 of the algorithm described in Appendix C over the bottom half + 4 instead of bottom half of $L_{\rm sum}$ eigenvalues to unmix them (although I haven’t included the plots that result from doing this here). , but the association of the second cliff with sum/diff Fourier families is clearly approximately correct.

### v3

@@ 4 Results for a MLP trained on modular arithmetic
Next, we study the alignment of top eNTK eigenvectors with the ground-truth Fourier features in a MLP trained on modular arithmetic (Gromov, 2023).
@@ 4.1 Review of the setup
In the experiments below, we report results for a particular training run at $p=29$, $n=512$ and $\alpha=0.7$. However, we find qualitatively similar results for different seeds and other values of $p$. (See Appendix F for a robustness check.)
With these hyperparameters, the model exhibits grokking (Power et al., 2022; Nanda et al., 2023). Namely, some time after reaching 100% accuracy on the training set, it suddenly goes from 0% to 100% accuracy on the test set (see the left panel of Fig. 3), suggesting that it switched to using a more generalizable representation of the underlying modular arithmetic algorithm.
Moreover, in this case we know the exact solution that the model learns after grokking (Gromov, 2023). At the end of training, the model solves modular addition by learning the weights
@@ 4.2 Results
Figure 2: Full and layerwise eNTK spectrum of the MLP trained on modular arithmetic after training to convergence (500 epochs). The left column displays the spectrum of the full eNTK, the middle column the spectrum of the layerwise eNTK wrt the parameters (9) in layer 1, and the right column the spectrum of the layerwise eNTK wrt the parameters (10) in layer 2.
@@
Time evolution of the spectrum. The second cliff is absent before grokking and appears at the grokking phase transition (see the right panel of Fig. 3). This suggests that feature-finding aside, perhaps the eNTK can be used to find phase transitions. In contrast to the eigenvalues, the eigenvectors evolve smoothly on either side of the phase transition. We comment on the evolution of the eigenvectors in Appendix C.
On the other hand, the first cliff is present already at initialization (see the r.h.s. of Fig. 3). This follows from the structure of the Layer 1 kernel, as we explain briefly in Appendix B, and suggests (in combination with the discussion below) that this model’s use of Fourier features could perhaps have been predicted at initialization from the eNTK.
Figure 3: Spectral change at the grokking phase transition for the MLP trained on modular arithmetic. Left: Train and test accuracy shows evidence for a grokking phase transition (sudden onset of generalization) around epoch 90. Right: at the same training time, the eNTK spectrum develops a kink at the base of the second cliff.
@@
Table 1: Frobenius norm of the eigendirections in each cliff with different Fourier feature families, as well as a control of normalized random vectors of the same dimension. A score of 28, which is the dimension of each Fourier feature family, would denote perfect alignment. This exercise shows strong alignment of the first cliff with the $a$ and $b$ Fourier feature families, and of the second cliff with the $a\pm b$ feature families.
Cliff a b sum diff ctrl
1 27.54 27.56 0.13 0.38 1.83
2 0.27 0.25 26.53 24.86 1.80
To check whether the span of each cliff contains a particular Fourier feature family, we compute the Frobenius norm between the cliff directions and the feature family, i.e. the sum of squared inner products between all pairs of directions of a normalized basis for the cliff and each feature family. The results, shown in Table 1, strongly suggest that the span of the first cliff contains the $a$ and $b$ Fourier families and the span of the second cliff contains the $a\pm b$ Fourier families.
Unlike in the TMS experiment however, the heat map of inner products of the eNTK eigenvectors in each cliff with feature vectors for the Fourier families is not 1:1. This degeneracy reflects the exact shift symmetry of the dataset under shifts in $a$ and $b$. Given knowledge from Table 1 of which Fourier families are contained within each cliff, we can surface the exact linear combinations of eNTK eigenvectors that align with each Fourier mode by ordering directions in each cliff eigenspace according to how smoothly they vary along the input axes in the directions specified in Table 1. We discuss this in Appendix C.
5 Results for a Transformer trained on modular arithmetic
In our third example, we study the alignment of eNTK eigenvectors with ground-truth Fourier features in a Transformer trained on modular arithmetic with cross-entropy loss (Nanda et al., 2023). This model learns a different ground-truth algorithm than the MLP trained on modular arithmetic that was the subject of section 4.
5.1 Review of the setup
In this experiment, for a fixed integer $p$ we construct a dataset consisting of the $p^{2}$ unique pairs of integers from $(0,\dots,p-1)$ represented as a sequence of three tokens [a,b,=]``[a,b,=]", labeled by their one-hot-encoded sum mod $p$ at the final token position. As in the previous experiment, we split the dataset into non-overlapping train and test sets controlled by a fractional hyperparameter $\alpha$. We then train a 1-layer GPT-type Transformer, without LayerNorm and with untied embedding/unembedding matrices, to learn the training set with cross-entropy loss and an AdamW optimizer. For a review of the exact architecture, see Appendix A of (Nanda et al., 2023).
Below, we report results for $p=29$, $d_{model}=64$, 4 attention heads of dimension $d/4=16$, and $n=256$ hidden units in the MLP. We use learning rate 3e-4, weight decay parameter 1.0, and $\alpha=0.5$. With these hyperparameters, the model exhibits grokking across several seeds.
The ground-truth modular arithmetic algorithm learned by this model after grokking is different from the one learned by the MLP in section 4 (Nanda et al., 2023). One major difference is that rather than learn Fourier modes at all intermediate frequencies, the model learns only Fourier modes at a sparse set of random, seed-dependent frequencies. Specifically, at the embedding step, the model maps the inputs $a$ and $b$ to sines and cosines at a set of key random frequencies $\omega_{k}=2\pi k/p$, $k\in\mathbb{N}$. See Fig. 4 for an example of key frequencies in a particular training run, obtained by applying a Fourier transform along the input dimension of the embedding matrix $W_{E}$ then computing the L2 norm in the other dimension. We will analyze this training run in the rest of this section. However, we observed similar phenomena at other random seeds. (See Appendix G for a robustness check.)
After the embedding step isolates the key frequencies, the model then computes the trigonometric sum identities for $\cos(\omega_{k}(a+b))$ and $\sin(\omega_{k}(a+b))$ on those frequencies only in the attention and MLP layers, and the trigonometric difference identity $\cos(\omega_{k}(a+b-c))$ on those frequencies only in the MLP output and unembedding layers. Finally, it uses the unembedding matrix to add together the $\cos(\omega_{k}(a+b-c))$’s for different $k$’s. This causes destructive interference unless a+bc=0a+b-c=0.
Figure 4: The L2 norms of the Fourier components of the embedding matrix $W_{E}$ in a training run of the Transformer model with the hyperparameters mentioned in the main text. Different seeds give different key frequencies. In this training run, there are two key frequencies at $k=1$ and 12, since modes with $k>\lfloor\frac{p}{2}\rfloor$ are redundant.
5.2 Results
Figure 5: Layerwise eNTK spectra of the Transformer trained on modular arithmetic wrt the “O” + “V” weight matrices in the attention block; the input matrix $W_{in}$ of the MLP layer; the output matrix $W_{out}$ of the MLP layer; and the unembedding matrix (from left to right, respectively), in a run with clear spectral separation. For additional seeds, see the discussion in Appendix G.
Leading layerwise eNTK eigenspaces align with Fourier features at key frequencies. Because we did not use LayerNorm in this experiment (in keeping with the conventions of (Nanda et al., 2023)), the eNTK spectra of different layers of this model differ from each other by orders of magnitude, and the full eNTK spectrum is dominated by the largest such contributions. It’s therefore more meaningful to look at the layerwise eNTK spectra.
In this run, we find that the leading five eigendirections of the layerwise eNTK associated with the O+V parameters in the attention block, the input matrix $W_{in}$ of the MLP layer, and the output matrix $W_{out}$ of the MLP layer each overlap significantly with the sum of $a$ and $b$ Fourier features $\{\cos(\omega_{k}a)+\cos(\omega_{k}b),\sin(\omega_{k}a)+\sin(\omega_{k}b)\}$ at the key frequencies shown in Fig. 4, while the top five eigendirections of the eNTK for the unembedding matrix overlap significantly with the ‘sum’ Fourier features. To see this, we compute the Frobenius norm between said eNTK eigendirections and the corresponding feature families at the key frequencies in Table 2. The number five is one more than twice the number of key frequencies in this run, suggesting that the eigendirections mix precisely those Fourier features with a zero mode or similar.  [^6] Across our earlier examples in sections 3 and 4, there was also a leading mode with a qualitatively large eigenvalue. We leave the question of why the leading direction here has a relatively smaller eigenvalue, so as to mix with the other top eigendirections, for future work. We also find that the next four eigendirections of the $W_{out}$ and unembedding layers overlap with the ‘sum’ and $a$ and $b$ Fourier feature families, respectively. In Appendix G, we show that this result generalizes across training runs with different (numbers of) key frequencies.
Intermittent emergence of sharp cliffs in the eNTK spectrum. In this run, the per-class averaged eNTK spectra at different layers are shown in Fig. 5. Each of these spectra contains a leading cliff ending at the fifth eigenvalue, while the $W_{out}$ and unembedding layers also contain secondary spectral cliffs of dimension 4. However, we do not find that sharp spectral gaps emerge consistently across other seeds trained at our fixed budget (see Appendix G). We leave the question of when sharp spectral gaps emerge in this model for future work.
Table 2: Frobenius norm of the eigendirections mentioned in the text with different Fourier feature families at the key frequencies, as well as a control of normalized random vectors of the same dimension. A score of 4 would denote perfect overlap with the Fourier feature directions.
layer A+B sum diff ctrl
OV 3.86 0.12 0.04 0.16
MLPin 3.89 0.05 0.01 0.14
MLPout (1) 3.90 0.02 0.03 0.14
MLPout (2) 0.02 3.94 0.01 0.14
U (1) 0.02 3.97 0.02 0.14
U (2) 3.92 0.00 0.02 0.10
@@ 6 Discussion
To summarize, we have provided evidence that eigenanalysis of the eNTK can surface learned features across three toy benchmarks commonly used by the mechanistic interpretability community.
For the toy models studied in this paper, it’s easy to write down closed-form expressions for the kernels in terms of the weights and data vectors. (See Appendices B and D.) It could be interesting to understand analytically what properties of said weights and data were needed to get the clean mechanistic signals presented here.
Another direction is to promote our checks that eNTK eigenvectors align with known ground-truth feature vectors to an approach that unsupervisedly matches a candidate eNTK eigendirection with a human-interpretable semantic meaning. To make progress on this problem, we can repurpose work on unsupervised feature discovery in the interpretability literature. In Appendix E, we show that for the modular arithmetic example in section 4, gradient ascent on a random input to maximize a scalar score built from an eNTK eigenvector that corresponds to a Fourier mode at a particular frequency yields synthetic inputs that resemble sinusoids of the same frequency. This provides a proof-of-concept example of how one could unsupervisedly assign semantic meanings to eNTK eigendirections.
Finally, we would ultimately like to make contact with realistic models, understanding if and how the eNTK can slot into an end-to-end pipeline for unsupervised feature discovery at scale. To this end, it would be interesting to compute and study the leading eNTK eigendirections  [^7] In larger models, we would presumably want to use approximation methods to compute the top eigendirections without ever materializing the full eNTK. in larger models, e.g. (Bricken et al., 2023), OthelloGPT (Li et al., 2023) and eventually realistic LLMs. We hope to report on this in the future.
Acknowledgments
I am grateful to Ari Brill, Tom Ingebretsen Carlson, Lauren Greenspan, Andrew Mack, Logan Smith, Lucas Teixeira and Dmitry Vaintrob for discussions and comments on the manuscript. This work was supported by PIBBSS and the Long-Term Future Fund.
@@ Appendix A Review of NTK theory
In this section, we review the derivation of Eq. (1), which motivates this line of study. As mentioned in the Introduction, a main claim of the NTK theory is that a neural network trained under gradient flow with MSE loss learns the function (1), (2) at the end of training as long as its trajectory stays in a “lazy limit” where the drift in parameters throughout training is parametrically small.
The idea behind the derivation is that if we take a limit where the changes in a model’s weights $W_{\mu}$ during training are small enough that we can truncate a Taylor expansion of the model in its weights to first order while maintaining a good approximation, then perhaps we can resum the linearized approximation to how the model changes at each step of gradient descent to get a closed-form formula for the function learned by the model.
@@
where the $\dots$’s denote terms quadratic and higher in $dW_{\mu}$. Using the definition of gradient descent to replace $dW_{\mu}\rightarrow-\eta\frac{d\mathcal{L}}{dW_{\mu}}$ for $\eta$ the learning rate and $\mathcal{L}$ the loss function, as well as the fact that the loss function itself depends on the model evaluated on points in the training set, we can massage (12) into
$f_{i}(x,W_{\mu}(t+1))-f_{i}(x,W_{\mu}(t))=\left.-\eta\sum_{j}\sum_{\alpha\in\mathcal{D}}\frac{d\mathcal{L}}{df_{j}(x_{\alpha})}\left[\sum_{\mu}\frac{df_{i}(x)}{dW_{\mu}}\frac{df_{j}(x_{\alpha})}{dW_{\mu}}\right]\right|_{W_{\mu}=W_{\mu}(t)}+\dots\,.$ (13)
The term in brackets on the RHS is the NTK, (2).  [^8] Throughout this section we also use that the NTK itself is time independent in the lazy limit.
Now suppose that we’re allowed to truncate the $\dots$ while maintaining a controlled approximation to the function learned by the neural network. Moreover, let’s specialize to MSE loss. This yields
@@
Appendix B Kernels used in the MLPs studied in this paper
@@
For the MLP/modular arithmetic experiment, $f(x)=W^{(2)}(W^{(1)}x)^{2}$ (6) implies the layer-1 Jacobian
@@
It could be interesting, and probably is tractable, to understand which properties of the weights and data lead to the mechanistic structure found in the main text. In general, we will leave this for future work.
As an appetizer however, note that an especially striking result in the main text - the fact that for the MLP trained on modular arithmetic, the Layer 1 eigenvalue cliff was present already at initialization, and hence the use of Fourier features could perhaps have been predicted at initialization with the eNTK - is easily explained by the structure of Eq. (18). Eq. (18) contains a parameter-agnostic factor of $x_{1}\cdot x_{2}$ which is zero for one-hot-encoded data points $(a,b)$ and $(a^{\prime},b^{\prime})$ unless either $a=a^{\prime}$ or $b=b^{\prime}$. This makes the Layer 1 kernel have nonzero values along each fixed-$a$ and fixed-$b$ line, and zeroes elsewhere. Moreover, the nonzero kernel entries at initialization are approximately identical for each choice of $a$ or $b$ by symmetry. Hence, the kernel at initialization has an approximate shift symmetry along the $a$ and $b$ axes, so is approximately diagonalized by Fourier waves on each axis, with (nonzero) eigenvalues related to the weight initializations.
In this example, the $x_{1}\cdot x_{2}$ factor in (18) relied on the model architecture (specifically, the fact that each one-hot-encoded data point only touches two columns of the Layer 1 weight matrix), and its conversion to a shift-symmetric kernel relied on the shift symmetry of the dataset. This illustrates a broader point that the NTK (and its eigenstructure) inherits information about both already at initialization.
Appendix C Algorithm to find a basis of cliff eigenvectors that line up with frequency-sorted Fourier features in the modular arithmetic experiment
In section 4, we saw by computing the Frobenius norm of the span of cliff eigenvectors with select Fourier feature families that the first cliff in the eNTK spectrum for a MLP trained on modular arithmetic contains the $a$ and $b$ Fourier families. However, the heat map of inner products of naive eNTK eigenvectors in the first cliff with feature vectors for the $a$ and $b$ Fourier families is not 1:1. See the left-hand side of Fig. 6, where in each square we plot the squared norm of the inner product of the $i$th eNTK eigenvector in the cliff with the $(\cos a,\sin a)$ or $(\cos b,\sin b)$ feature families.  [^9] This equals the maximal correlation with a phase-shifted cosine in (11), i.e. the best match over possible phases $\varphi_{k}^{(1)}$, $\varphi_{k}^{(2)}$. This is because the cliff eigenvalues are almost exactly degenerate due to the exact shift symmetry of the dataset, so the eigenvectors are only defined up to a random rotation.
To find an orthogonal basis within the Layer 1 cliff that lines up 1:1 with the Fourier features, we can apply the following two-stage “graph smoothness” algorithm:
Figure 6: Results from applying the two-stage graph smoothness algorithm to rotate the eNTK eigenvector basis in the Layer 1 cliff. Left: heatmap showing the squared norm of the inner product of naive eNTK eigenvectors in the cliff with the $(\cos a,\sin a)$ or $(\cos b,\sin b)$ feature families. Right: heatmap showing the squared norm of the inner product of basis vectors in the Layer 1 cliff found from applying the two-stage graph smoothness algorithm with the $(\cos a,\sin a)$ or $(\cos b,\sin b)$ feature families.
@@
    penalizing cases where entries in the (length $p^{2}$) vector $v$ corresponding to lattice points that differ by 1 in the $a,b$ directions respectively are not the same. In practice, the matrices $L_{a}$, $L_{b}$ that realize (21), (22) are $L_{1}\otimes I_{p}$ and $I_{p}\otimes L_{1}$ respectively, for $L_{1}$ the dimension $(p,p)$ matrix with 2 on the diagonal and 1-1 along diagonals offset by 1 and at the wrap-around entries $(1,p)$ and $(p,1)$.
@@
    We then diagonalize one of the two Laplacians, WLOG $L_{a}$, inside the span of the Layer 1 cliff. Concretely, we first compress $L_{a}$ by computing $A=c^{T}L_{a}c$ for $c\in\mathbb{R}^{N,k}$ the column-orthonormalized collection of eNTK eigenvectors in the Layer 1 cliff, then eigendecompose the compressed Laplacian finding $A=U\Sigma U^{T}$ with $U\in\mathbb{R}^{k,k}$, and finally rotate the eNTK cliff by the eigenvector matrix of the compressed Laplacian, $c^{\prime}=cU$. This sorts the directions in the cliff by $a$-smoothness.  [^10] This is because directions where we minimize $v^{T}L_{a}v$ in a given vector space are the smoothest directions in the vector space by the operational definition (21) of the Laplacian, and we can minimize it from an eigendecomposition by the Rayleigh-Ritz theorem.
@@
    Finally, we take the directions in the cliff corresponding to the bottom half of the $L_{a}$ eigenvalues  [^11] I.e., the bottom $2\lfloor\frac{p}{2}\rfloor$ directions in practice. and diagonalize $L_{b}$ inside that span, using the same procedure as in step 2. This sorts those directions by $b$-smoothness.
The logic is that if the Layer 1 cliff was spanned exactly by the $(\cos a,\sin a),(\cos b,\sin b)$ Fourier families, then we would expect the spectrum wrt $L_{a}$ to contain degenerate near-zero values corresponding to the ‘$b$’ Fourier modes that are constant in the ‘$a$’ direction, followed by a tower of increasing pairs of nonzero eigenvalues corresponding to the sorted ‘$a$’ Fourier modes.
Applying this algorithm to the span of the Layer 1 cliff, we indeed find a set of basis directions in 1:1 correspondence with frequency-sorted $a$ and $b$ Fourier modes. This is shown on the right-hand side of Fig. 6.
An analogous rotation works to surface the ‘sum’ and ‘difference’ Fourier feature vectors as linear combinations of eigenvectors in the second eNTK cliff (see the left-hand column of Fig. 2). In this case, we apply the two-stage graph smoothness algorithm to the sum and difference directions, first constructing Laplacians
@@
that penalize variation along the diagonals, then diagonalizing first the sum then difference Laplacian within the span of the second eigenvalue cliff. This situation is equivalent to two-stage diagonalization in the $a$ and $b$ directions up to a 45-degree rotation. In particular, $(a,b)$ and $(a+1,b+1)$ have the same aba-b, so the difference family is flat wrt $L_{\rm sum}$, and vice versa.
@@
Figure 7: Results from applying the two-stage graph smoothness algorithm wrt the sum and difference Laplacians (23), (24) to rotate the eNTK eigenvectors in the second cliff of the MLP/modular arithmetic experiment in section 4 at different epochs. Each plot is a heatmap showing the squared norm of the inner product of rotated basis vectors in the second cliff with the $(\cos a+b,\sin a+b)$ or $(\cos a-b,\sin a-b)$ feature families.
@@
The results are shown in Fig. 7. From the bottom right plot showing the results of the rotation at the end of training, we see that the procedure indeed surfaces a basis in the second eigenvalue cliff that (approximately) lines up with the sum and difference Fourier features. The heatmaps for each family form a “>>” shape because stepping across the diagonal advances the argument of each Fourier mode by $2\cdot 2\pi k/p$, so the Laplacian energy is minimal near $k\approx 0$ and $k\approx p/2$, and maximal near $k\approx p/4$.  [^12] Note that this rotation is imperfect because the low-energy k=1``k=1" and k=14``k=14" sum modes mix with the $L_{\rm diff}$ modes. Indeed, one can make the heatmaps in Fig 7 appear closer to 1:1 by diagonalizing $L_{\rm diff}$ over the bottom half + 4 instead of bottom half of $L_{\rm sum}$ eigenvalues to unmix them, although we haven’t included the plots that result from doing this here.